OpenStudy (anonymous):

Shifting parent graphs to Graph Equations: y=-2(x-3)^2+4

OpenStudy (anonymous):

I would assume you are to expand this and simplify y = -2(x - 3)(x - 3) + 4 y = -2(x^2 - 6x + 9) + 4 y = -2x^2 + 12x - 18 + 4 y = -2x^2 + 12x - 14

OpenStudy (amistre64):

the (-) slope tells us its upside down the (x-3) tells us we moved it 3 place to the right ()^2 tells us its a parabola the +4 on the end says we moved it up 4 places

OpenStudy (anonymous):

Okay but how do i simplify it

OpenStudy (amistre64):

its already simplified in the form you gave us

OpenStudy (anonymous):

okay so the question that i have to answer is how do i put it in a graphing equation

OpenStudy (amistre64):

like a graphing calcultor equation?

OpenStudy (amistre64):

otherwise, it its got a y = something with an x, its good to graph :)